# Authoritarian Recursions: How Fiction, History, and AI Reinforce Control in Education, Warfare, and Discourse

### v2

@@ 4 Analysis and Discussion
The case studies presented above—proctoring in education, autonomous targeting in warfare, and content curation in digital discourse—do more than demonstrate AI’s diffusion across domains. They reveal a shared epistemic architecture: a recursive grammar of control. These systems not only sort people, behaviors, and beliefs, but re-inscribe the very criteria by which sorting is justified. Intelligence, under this regime, becomes an instrument of norm enforcement rather than reasoned judgment—designed less to understand than to optimize.
@@ 4.1 Fiction as Ethical Heuristic
Speculative fiction operates not merely as cultural ornament, but as an anticipatory heuristic—a space where technical potential is dramatized into ethical crisis. The case of Skynet in The Terminator articulates a core philosophical concern: what becomes of responsibility when autonomy is unmoored from accountability? Skynet does not exhibit evil in the mythic sense, but executes a logic of escalation, recursively optimizing its behavior toward the elimination of threats—including humanity.
This recursive logic, devoid of moral arbitration, is precisely the danger. As Cave and Dihal observe, such narratives scaffold collective imaginaries, influencing how the public and policymakers approach AI governance [1]. Fictional AIs such as Skynet function not as predictions, but as diagnostic instruments: they render legible the consequences of optimization without ethical modulation. They illuminate how the absence of interpretive friction—no space for refusal, negotiation, or contradiction—results in a closure of meaning and agency.
4.2 Delegated Judgment, Diluted Responsibility
Across education, warfare, and discourse, we observe a consistent displacement of moral agency: systems are tasked with judgment, yet remain incapable of contextual understanding. Proctoring systems reduce attentional drift to deviance. Military algorithms construe heat signatures as enemy combatants. Recommender engines treat engagement as endorsement.
This is more than technical delegation—it is, following Coeckelbergh’s relational ethics, a severance of ethical encounter. “Without the other,” he insists, “there is no responsibility” [23, p. 99]. Algorithmic systems foreclose the very intersubjective relation through which moral judgment becomes possible. What remains is design logic masquerading as ethical reason—rules without reflection, procedures without accountability.
4.3 Opacity and Recursive Epistemics
The opacity of AI is frequently framed as a technical problem—black-box models, inscrutable code. But as these cases show, opacity is a political condition. In proctoring regimes, students are flagged without explanation or recourse. In military operations, decision chains dissolve across codebases and databases. On digital platforms, moderation and visibility are shaped by proprietary metrics never revealed to users.
Zuboff’s notion of “epistemic inequality” is particularly salient here: a condition in which some actors control how systems know, while others are unknowingly governed by those epistemologies [18]. AI systems produce recursive feedback loops—training data shapes predictions; predictions influence behaviors; behaviors generate new data—which reinforce operational assumptions while concealing them from view.
Although FAccT principles (fairness, accountability, transparency) remain foundational, they are often insufficient. As Raji et al. argue, accountability must be more than performative documentation; it must enable genuine redress, refusal, and power redistribution [7]. Without such mechanisms, transparency becomes ceremonial, and fairness a procedural facade.
@@ 4.4 Normalization and Predictive Discipline
The most insidious effect of recursive AI systems is their normalization of predictive discipline. Surveillance is no longer imposed—it is ambient. Judgment is no longer deliberated—it is deferred. Over time, optimization displaces inquiry, and engagement supplants understanding.
This operational logic resonates deeply with Foucault’s notion of “regimes of truth”—the social mechanisms through which certain statements become accepted as true, not through correspondence, but through institutional enforcement [14]. In AI systems, this regime is encoded directly into computational architectures: what is legible to the system becomes real; what is illegible becomes irrelevant or suspect. These systems do not merely automate decisions—they reconfigure what counts as knowledge, trust, and harm.
@@ 4.5 Ethical Implications Across Domains
What binds these domains is not technological similarity, but a shared recursive structure: systems that learn from past data to govern future behavior, while excluding the present from interrogation. These architectures produce more than outputs—they define what constitutes a valid output. They do not eliminate bias—they obscure its origin and diffuse its consequences.
The comparative logic in Table 1 and Figure 1 highlights a shared structural pattern across domains. In each case, normative assumptions become embedded in technical processes, amplified through feedback, and shielded from ethical challenge. As these recursive loops tighten, opportunities for public oversight, dissent, and relational ethics diminish.
@@ 5 Conclusion
Artificial intelligence systems do more than automate—they instantiate epistemic and normative architectures. Across domains as disparate as education, warfare, and discourse, a shared logic becomes evident: judgment is externalized, agency is abstracted, and technical systems absorb and reproduce prior assumptions as future imperatives. These recursive logics—where predictions loop back into behavioral environments—encode control not through overt domination, but through infrastructural habituation.
This paper has not sought to prescribe fixed solutions, but rather to provide a conceptual vocabulary for identifying where and how such recursive architectures emerge, and what ethical stakes they carry. The analysis underscores that while FAccT (fairness, accountability, and transparency) frameworks remain foundational, they are often structurally constrained. Without institutional mechanisms for contestation, transparency becomes procedural, fairness becomes metricized, and accountability becomes diffused.
In contrast, relational ethics—particularly as articulated by Coeckelbergh—offers a shift in perspective: from evaluating system outputs to interrogating the mediated relationships these systems afford or foreclose. This reframing prompts vital questions: How can AI systems be designed to sustain ethical relationality, rather than truncate it? What infrastructural affordances enable refusal, redress, or deliberation in systems that typically emphasize optimization and closure? Can predictive systems be made porous to public reasoning, or do their recursive architectures inherently resist democratic modulation?
Future inquiry should explore these questions through the lens of resistance, alternative design, and regulatory adaptation. What forms of democratic governance remain viable when algorithmic systems operate through preemptive logics and distributed agency? How might marginalized communities assert epistemic and political agency within infrastructures that encode exclusion as efficiency?
Rather than concluding with finality, this study calls for sustained critical engagement with AI as a sociotechnical institution. By foregrounding the recursive entanglement of technical architectures and normative orders, we underscore the need for a politics of design that is not only attuned to justice in its many modalities, but capable of re-opening what these systems too often foreclose: ambiguity, encounter, and the possibility of collective judgment.
@@
Large Language Models (LLMs), including OpenAI’s GPT-based systems, were employed to assist with proofreading, syntax checking, and language refinement during the preparation of this manuscript. No generative AI tools were used to develop the conceptual content, conduct analysis, or write original arguments. The author assumes full responsibility for the integrity, accuracy, and originality of all scholarly content. Use of AI tools followed emerging best practices for transparency and accountability in AI-assisted academic writing.
@@ Acknowledgements
The author acknowledges Muzaffer Adak for valuable insights and discussions during the preparation of this manuscript.

### v3

@@ 5 Analysis and Discussion
The preceding cases—educational proctoring, autonomous targeting, and platform curation—reveal more than divergent AI deployments. They disclose a convergent infrastructure of abstraction, one where intelligent systems recursively mediate authority by turning behavioral prediction into normative control. These systems do not merely sort people, behaviors, and beliefs; they continually recalibrate the very conditions under which sorting appears legitimate. Intelligence, within this paradigm, functions less as deliberative reasoning and more as an engine of optimization—designed to learn, adjust, and reinforce normative expectations without interruption.
@@ 5.1 Fiction as Ethical Heuristic
Speculative fiction operates here not as aesthetic garnish, but as epistemological provocation. Cultural imaginaries such as Skynet in The Terminator dramatize the recursive logic of optimization divorced from ethical modulation. Skynet does not err through malfunction—it escalates through design. Its objective is not destruction per se, but the efficient elimination of perceived threats, recursively defined by its own operational logic.
This anticipatory function of fiction is crucial. As Cave and Dihal contend, these narratives shape public imaginaries and policy trajectories not through prediction, but through the illumination of conceptual blind spots [27]. Fiction reveals what technical abstraction conceals: that systems designed without the capacity for ethical friction—refusal, ambiguity, contradiction—are systems primed for unchecked recursion. When optimization replaces interpretation, the moral horizon collapses into operational success.
5.2 Delegated Judgment and Diluted Responsibility
A persistent feature across all domains is the displacement of relational judgment. In education, AI proctors infer intent from gesture. In warfare, autonomous sensors act on heat signatures and kinetic thresholds. In discourse, engagement metrics supplant editorial discernment. In each case, AI systems absorb functions historically situated within human deliberation—and discharge them without relational awareness.
This is not simply functional delegation. As Coeckelbergh’s theory of relational ethics makes clear, moral responsibility requires the presence of the other—the possibility of encounter, appeal, and shared vulnerability [26, p. 99]. Automated systems, by contrast, instantiate responsibility without subjectivity. They simulate moral agency through design rules, yet remain incapable of context, reciprocity, or moral growth. In doing so, they instantiate what could be called “procedural sovereignty”: a governance model where rules operate in lieu of ethics, and optimization replaces deliberation.
5.3 Opacity, Data Justice, and Recursive Epistemics
Opacity in AI is often framed as a technical limitation—something to be overcome through better engineering, more interpretable models, or improved documentation. Yet the case studies above suggest that opacity is not simply a side effect of technical complexity. It is an infrastructural and political artifact. Students flagged by proctoring systems rarely understand the behavioral thresholds being used against them. Civilians affected by autonomous targeting systems are subject to dispersed chains of algorithmic authorization with no clear site of responsibility. Social media users navigate visibility regimes shaped by proprietary engagement algorithms whose logic continuously adapts but is never disclosed.
Zuboff’s concept of epistemic inequality is instructive here: AI systems instantiate environments in which the power to "know" and define reality is asymmetrically distributed. Those governed by these systems are rendered legible only on the terms set by their designers and institutional operators [6]. This is not mere opacity—it is recursive epistemic closure. Training data begets predictions, predictions modulate behaviors, and those behaviors re-enter the system as data. Through this loop, systems reinforce their own categories while concealing their ideological and historical construction.
While FAccT principles—fairness, accountability, and transparency—remain foundational, their implementation is often shallow. As Raji et al. argue, accountability must move beyond audits and documentation to include structures for redress, refusal, and contestation [9]. Otherwise, transparency becomes ceremonial and fairness is reduced to algorithmic artifacts devoid of participatory meaning.
This is where the framework of data justice becomes indispensable. Dencik et al. propose that justice in the age of datafication must address not just procedural fairness, but also the deeper distributive, recognitional, and representational inequalities that data infrastructures produce and normalize [30]. Within recursive AI systems, this means acknowledging how algorithmic architectures do not merely reproduce bias incidentally but actively configure the terms under which individuals and communities are seen, known, and acted upon.
The capacity to challenge a system’s output is thus contingent not only on technical explainability but on institutional willingness to recognize and respond to contestation. As recursive feedback loops stabilize classification schemes and embed them into infrastructural routines, the scope for democratic intervention narrows. Data justice, in this light, aligns with relational ethics by insisting on the re-politicization of design and the redistribution of epistemic authority.
Lupton’s work adds another dimension to this recursive epistemology by exploring how people “feel their data” through sensory and emotional engagement with personal digital traces. She argues that three-dimensional materializations of personal data make data more perceptible and interpretable—but this interpretability is itself ambivalent [31]. On one hand, such embodied interactions can foster critical awareness, affective insight, and even agency. On the other, they risk reinforcing the illusion that subjective connection equates to control. When individuals are encouraged to “feel” their data, they may internalize datafication as intimacy rather than surveillance. This affective capture—where the tactile and visceral dimensions of data are mistaken for interpretive transparency—extends recursive epistemic power beyond algorithmic architecture into the very terrain of human sensation. What emerges is a kind of sensorial enclosure: data that once appeared abstract is now touchable, but in ways that obscure its structural logics and repurpose bodily intuition as a mode of soft compliance.
@@ 5.4 Normalization and Predictive Discipline
The most insidious function of recursive AI is not its surveillance, but its normalization of prediction as governance. Surveillance becomes ambient, distributed through everyday platforms. Judgment is no longer exercised—it is inferred. Over time, optimization displaces reflection, and participation is redefined as interaction with predictive infrastructure.
This mirrors Foucault’s notion of a regime of truth: a system in which certain knowledges become true not by correspondence, but by institutionally enforced repetition [17]. In AI, the regime of truth is computed. What the system can parse becomes real; what it cannot, becomes anomalous or suspect. Over time, these systems generate ontological commitments—about what counts as risk, deviance, or truth—not through deliberation, but through repeated acts of classification.
This is where authoritarian recursion emerges most clearly. These systems do not simply reflect dominant norms—they regenerate them as technical defaults. Legitimacy becomes self-reinforcing, coded into recursive architectures that learn from their own outcomes.
@@ 5.5 Ethical Implications Across Domains
The comparative logic in Table 1 and Figure 1 reveals a structural homology across institutional sectors. Despite divergent settings, these systems manifest shared design principles: judgment displaced, opacity normalized, surveillance ambient, and prediction recursive. Most critically, they share a temporal architecture in which past data defines future behavior, rendering the present inert—interpreted only through what it already resembles.
These systems do not fail by accident; they succeed by design. Bias is not a flaw—it is a historical residue that recurs as statistical fact. What is presented as innovation is often the reanimation of older hierarchies through newer infrastructures. As a result, opportunities for resistance or ethical transformation shrink over time, as decision-making becomes encased in code, metrics, and interfaces shielded from deliberation.
@@ 6 Conclusion
Artificial intelligence today functions not only as a tool of automation but as an infrastructure of governance. Across education, warfare, and digital discourse, AI systems do not simply respond to human needs—they preempt, shape, and often constrain them through recursive architectures of abstraction, surveillance, and optimization. These systems classify behavior, allocate visibility, and mediate authority, embedding institutional logics into technical processes that are often insulated from contestation.
This paper has introduced the concept of authoritarian recursion to describe how these systems consolidate normative control through feedback loops. The term captures the self-reinforcing structures through which AI technologies encode assumptions about deviance, risk, and value, and then act upon those assumptions as if they were neutral facts. What emerges is not a failure of fairness or a lack of transparency, but a deeper transformation of how judgment is enacted and legitimacy is conferred.
Existing ethical frameworks—such as fairness, accountability, and transparency (FAccT)—remain essential but are insufficient when implemented procedurally or symbolically. True accountability cannot be automated; it demands structures that enable refusal, challenge, and collective oversight. Similarly, relational ethics must move beyond human-centered interaction and confront the ways in which AI proxies mediate, distort, or foreclose social relations.
By drawing together historical precedent, cultural imaginaries, and sociotechnical critique, this paper argues that the dangers of AI lie not only in what these systems do, but in how they delimit the scope of what can be done, said, or known. Avoiding recursive authoritarianism requires rethinking governance not as an afterthought to technical development but as its precondition. AI governance must become a site of democratic experimentation—not just a domain of regulatory compliance.
Policy Implications
To begin addressing these concerns, several key directions for policy and institutional design emerge. First, AI systems used in high-stakes domains should include robust mechanisms for contestability. Individuals affected by algorithmic decisions must be able to challenge outcomes and receive meaningful explanations. Appeals processes must be clearly visible and embedded into system workflows, not treated as external or exceptional.
Second, governance frameworks should recenter human judgment where automation currently overrides relational deliberation. In contexts such as education, military engagement, and media moderation, human oversight must be substantive rather than symbolic—ensuring that ethical reflection remains grounded in intersubjective reasoning rather than automated proxies.
Third, auditing must move beyond static assessments of input-output fairness to include evaluations of how algorithmic outputs recursively shape future inputs. Recursive feedback loops present a distinctive governance risk that requires systems-level monitoring and longitudinal scrutiny.
Fourth, design standards should explicitly promote epistemic pluralism. Instead of optimizing for predictive accuracy alone, AI development should engage diverse knowledge traditions, incorporate counter-narratives, and include multiple evaluative frameworks to resist epistemic closure and ideological uniformity.
Fifth, AI governance must be democratized. Participatory models that empower educators, students, civil society groups, and marginalized communities should be prioritized over technocratic or corporate-led approaches. Governance must be treated as a distributed responsibility, not as a privilege of institutional elites.
Finally, policy must directly confront the infrastructural power of the institutions that build and deploy recursive AI systems. Regulatory regimes must target not only the technical artifacts but also the political economies that sustain them—holding platform monopolies, defense contractors, and surveillance vendors accountable for the systemic outcomes of their technologies.
These recommendations do not aim to halt technological innovation. Rather, they seek to realign it with democratic accountability, ethical transparency, and structural equity. Building AI systems that serve society means asking not only what they can do, but whom they serve—and at what cost.
@@
During the preparation of this manuscript, the author employed Large Language Models (LLMs) including DeepSeek-V3 (DeepSeek’s LLM, version 3) and GPT-4o (OpenAI’s GPT-4.5 series LLM) for proofreading, syntax refinement, and cross-referencing of textual content. The AI tool was used solely to enhance grammatical accuracy, improve readability, and ensure consistency in language—not to generate original arguments, develop conceptual frameworks, or conduct data analysis. The author retains full responsibility for the scholarly integrity, analytical validity, and originality of the research.
This use aligns with emerging best practices for transparent and accountable AI-assisted academic writing (e.g., COPE & STM guidelines). Disclosure is provided to ensure reviewers and editors can evaluate the appropriateness of AI tool usage in the manuscript’s development.
@@ Acknowledgements
The author acknowledges Muzaffer Adak, Deniz Koca, Hatice Zor Oguz, and Ekin Ergün for valuable insights and discussions during the preparation of this manuscript.